OpenStudy (anonymous):

One square millimeter equals _____ square centimeter(s). (1 cm = 10 mm)

sam (.sam.):

\[1 cm = 10 mm\] \[\Rightarrow1^2 cm^2 = 10^2 mm^2\] \[\Rightarrow1cm^2=10^2mm^2\]

sam (.sam.):

\[1cm^2=100mm^2\] \[\frac{1}{100}cm^2=1mm^2\]

sam (.sam.):

Do you get it?

OpenStudy (anonymous):

the choices are 10^-1 , 10^-2 , 10^3, 10^-3 , but i dont understand lol

sam (.sam.):

Yeah from here \[\frac{1}{100}cm^2=1mm^2\] What is \(\frac{1}{100}\)?

sam (.sam.):



sam (.sam.):



sam (.sam.):

You have \(0.01cm^2=1mm^2\), they want the answer in scientific notation. What is 0.01 in scientific notation?